# Generative Adversarial Networks for High-Dimensional Item Factor Analysis: A Deep Adversarial Learning Algorithm

### v1

@@ Abstract
Advances in deep learning and representation learning have transformed item factor analysis (IFA) in the item response theory (IRT) literature by enabling more efficient and accurate parameter estimation. Variational Autoencoders (VAEs) have been one of the most impactful techniques in modeling high-dimensional latent variables in this context. However, the limited expressiveness of the inference model based on traditional VAEs can still hinder the estimation performance. This study introduces Adversarial Variational Bayes (AVB) algorithms as an improvement to VAEs for IFA with improved flexibility and accuracy. By bridging the strengths of VAEs and Generative Adversarial Networks (GANs), AVB incorporates an auxiliary discriminator network to reframe the estimation process as a two-player adversarial game and removes the restrictive assumption of standard normal distributions in the inference model. Theoretically, AVB can achieve similar or higher likelihood compared to VAEs. A further enhanced algorithm, Importance-weighted Adversarial Variational Bayes (IWAVB) is proposed and compared with Importance-weighted Autoencoders (IWAE). In an exploratory analysis of real empirical data, IWAVB demonstrated superior expressiveness by achieving a higher likelihood compared to IWAE. In confirmatory studies with simulated data, IWAVB achieved similar mean-square error results to IWAE while consistently achieving higher likelihoods. Moreover, in simulations where latent variables followed a multimodal distribution, IWAVB outperformed IWAE by providing more accurate parameter estimates. With its innovative use of GANs, IWAVB is shown to have the potential to extend IFA to handle large-scale data, facilitating the potential integration of psychometrics and multimodal data analysis.
@@ 1 Introduction
Psychological and educational research now frequently involves diverse data types, including large-scale assessment, high-volume data, and process data involving unstructured and textual responses (e.g., Weston and Laursen,, 2015; Zhang et al.,, 2023; Ma et al.,, 2024). Traditional estimation algorithms often struggle to efficiently process and analyze these data, leading to unstable model estimates, long computation times, and excessive memory usage. Deep Learning (DL) literature has much to offer to handle the above challenges. If we view IFA from the lens of generalized latent variable models (Muthén,, 2002; Rabe-Hesketh and Skrondal,, 2008), representation learning (Bengio et al.,, 2013) can help model and capture intricate latent patterns from observed data. Unlike traditional IFA estimation methods, which often relies on strict parametric and distributional assumptions, representation learning leverages powerful learning algorithms and expressive models such as neural networks to generate richer and highly flexible representations when modeling data. This improves the modeling and detection of subtle patterns for high-dimensional latent variables that classical methods might underfit. Therefore, by integrating representation learning with IFA, researchers can achieve a more comprehensive understanding of latent traits.
@@
An alternative representation paradigm, Generative Adversarial Networks (GANs, Goodfellow et al.,, 2014), a celebrated framework for estimating generative models in deep learning and computer vision, offers greater flexibility in modeling arbitrary probability densities and can contribute to the expressiveness. Multiple variants of GANs have achieved remarkable success in complex tasks such as image generation (e.g., Radford,, 2015; Arjovsky et al.,, 2017; Zhu et al., 2017a, ). Among these models, the combination of GANs and VAEs can be traced back to Adversarial Autoencoders (AAEs; Makhzani et al.,, 2015) and the VAE-GAN model (Larsen et al.,, 2016). Figure 1 provides an intuitive visual example demonstrating how integrating GANs and VAEs can lead to better representation learning. The figure compares ten handwritten images sampled from four models: (a) GAN, (b) WGAN (a GAN variant, Arjovsky et al.,, 2017), (c) VAE, and (d) VAE-GAN. While GAN generated only the digit 1, WGAN exhibited unstable performance, and the images produced by VAE appeared blurry. In contrast, VAE-GAN generated relatively clear results. However, vanilla GANs, VAE-GANs, and AAEs rely solely on adversarial loss, formulated as a minimax game between networks to deceive a discriminator. As a result, they are not motivated from a maximum-likelihood perspective and are therefore not directly applicable for improving marginal maximum likelihood estimation in IFA.
@@ 2 Adversarial Variational Bayes for Item Factor Analysis
In the context of IFA, given the observed responses $\bm{x}$, the Variational Inference (VI) method uses an inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ to approximate the posterior distribution of the latent variable $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$. The $\bm{\phi}$ represents the parameters for the inference model, and ${\bm{\theta}}$ includes the parameters for the posterior distributions (e.g., items’ parameters). As a specific approach to perform VI, Adversarial Variational Bayes can be regraded as an improved version of Variational Autoencoder which construct more expressive inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$, and is shown to improve the likelihood estimation of parameters for the Graded Response Model in this section.
@@ 2.1 The Graded Response Model
The Graded Response Model (GRM; Samejima,, 1969) is a specific model for IFA used for analyzing ordinal response data, particularly in psychological and educational measurement. The GRM assumes responses to an item are ordered categories (e.g., Likert scale). We follow the notation as in Cai, (2010) for this section.
Given $i=1,\dots,N$ distinct respondents and $j=1,\dots,M$ items, the response for respondent $i$ to item $j$ in $C_{j}$ graded categories is $x_{i,j}\in\left\{0,1,\dots,C_{j}-1\right\}$. Assuming the number of latent variables is $P$, the $P\times 1$ vector of latent variables for respondent $i$ is denoted as $\bm{z}_{i}$. For item $j$, the $P\times 1$ vector of loadings is $\bm{\beta}_{j}$, and the $(C_{j}-1)\times 1$ vector of strictly ordered category intercepts is $\bm{\alpha}_{j}=(\alpha_{j,1},\dots,\alpha_{j,C_{j}-1})^{T}$. Also, the parameters of item $j$ is denoted as $\bm{\theta}_{j}=(\bm{\alpha}_{j},\bm{\beta}_{j})$. The Graded Response Model is defined as the boundary probability conditional on the item parameters $\bm{\theta}_{j}$ and latent variables $\bm{z}_{i}$: for $k\in\left\{1,\dots,C_{j}-1\right\}$,
@@
The conditional probability for a specific response $x_{i,j}=k\in\left\{1,\dots,C_{j}-1\right\}$ is
@@ 2.2 Marginal Maximum Likelihood
For the respondent $i$, and the item $j$, with the Equation (2), the response $x_{i,j}$ follows the multinomial distribution with 1 trail and $C_{j}$ mutually exclusive outcomes. The probability mass function is:
$p_{\bm{\theta_{j}}}(x_{i,j}\mid\bm{z}_{i})=\prod\limits_{k=0}^{C_{j}-1}p_{i,j,k}^{\mathbbm{1}_{k}(x_{i,j})},$ (3)
where $\mathbbm{1}_{k}(x_{i,j})=1$ if $x_{i,j}=k$ and otherwise it is 0.
Considering all the responses $\bm{x}_{i}=(x_{i,j})_{j=1}^{M}$ of the respondent $i$, since each response is conditionally independent given the latent variables, the conditional probability is:
$p_{\bm{\theta}}(\bm{x}_{i}\mid\bm{z}_{i})=\prod\limits_{j=1}^{M}p_{\bm{\theta_{j}}}(x_{i,j}\mid\bm{z}_{i}),$ (4)
where $\bm{\theta}=(\bm{\theta}_{j})_{j=1}^{M}$.
Assuming the prior distribution of latent variables is $p(\bm{z}_{i})$ and given $p_{\bm{\theta}}(\bm{x}_{i}\mid\bm{z}_{i})$, the marginal likelihood of $\bm{x}_{i}$ is:
@@
Since the above equation contains $N$ integrals in high-dimensional latent space $\mathbb{R}^{P}$, it is computationally intensive to directly maximize the likelihood. Instead, variational methods that the current study focuses on improving, provide an approximation of $\log\mathcal{L}(\bm{\theta}\mid\bm{X})$ by one lower bound, which achieve efficient computation.
@@ 2.3 Variational Inference and Variational Autoencoder
For each set of item parameters $\bm{\theta}$, if there exists $\bm{\phi}$ such that $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ exactly match $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$ the KL divergence term will be zero, so the marginal maximum likelihood estimator $\bm{\theta}^{*}$ will also be the maximizer of ELBO. However, this will be a very ideal case. Generally speaking, for fixed $\bm{\theta}$, if ELBO is increased by taking gradient steps in $\bm{\phi}$ space, the KL divergence will more approach to its minimum. On the other hand, if $\bm{\phi}$ is fixed, increasing ELBO with respect to $\bm{\theta}$ will push the marginal likelihood $p_{\bm{\theta}}(\bm{x})$ higher.
Therefore, our objective is
@@ 2.4 Adversarial Variational Bayes
Therefore, we need to do two optimization for Objectives (12) and (14), summarized in Algorithm 1. Note that $T(\bm{x},\bm{z})$ and $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ can be implemented as neural networks. However, in the context of the GRM, the decoder $p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)$ corresponds directly to the likelihood definition for the GRM. As such, it is modeled by a single-layer neural network consisting of a linear function and a sigmoid activation.
@@ 2.5 Our Contributions to the Literature
Given that VAEs have been a powerful approach to perform exploratory IFA (Urban and Bauer,, 2021), the innovative contribution of the AVB method is to bring richer inference model choices for the implementation by deep learning. Over the past decade, deep learning techniques have thrived (LeCun et al.,, 2015), and Item Factor Analysis has derived considerable benefits from these advances. When applying neural networks to estimate parameters in IFA, key distinctions arise in the design of the network architecture and the choice of loss function. A straightforward strategy is to feed one-hot encodings [^1] In one-hot encoding, each respondent or item is represented by a binary vector whose length matches the total number of respondents or items. Exactly one element of this vector is set to 1 (indicating the specific respondent or item), and all other elements are set to 0. for both respondents and items into two separate feedforward neural networks: one estimates the latent variables of the respondents, while the other infers the item parameters. This approach typically employs a variation of the Joint Maximum Likelihood (JML) for its loss function, relying on the product of conditional probabilities in Equation (4) across all respondents (Tsutsumi et al.,, 2021; Zhang and Chen,, 2024). However, as demonstrated by Ghosh, (1995), the JML solution leads to inconsistent estimates for item parameters.
An alternative that proves more statistically robust relies on the marginal maximum likelihood procedure discussed earlier (Bock and Aitkin,, 1981; Wirth and Edwards,, 2007). Drawing on these perspectives, researchers merge representation learning for latent-variable models with neural architectures by adopting autoencoder designs. Through variational methods, they approximate maximum likelihood estimation for latent variable models. A notable example in the literature is the Importance-Weighted Autoencoder Estimation algorithm (Urban and Bauer,, 2021). In our study, we propose an importance-weighted adversarial variational Bayes, refines this approach.
@@
To address this limitation, the AVB framework modifies $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ from Equation (11) into a fully black-box inference mechanism that cooperates with adversarial objectives. Through this combined process, the algorithm derives an item parameter estimate $\bm{\theta}^{*}$ that optimizes the likelihood of the observed data, while also constructing a distribution $q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})$ that closely approximates the true posterior $p_{\bm{\theta}^{*}}(\bm{z}\mid\bm{x})$.
@@
Under misspecification such as inappropriate assumptions of latent variable distributions, AVB’s architecture can deliver a higher marginal likelihood theoretically. Meanwhile, by improving the marginal log-likelihood, one can reduce the reconstruction error $-\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]$, which measures the disparity between real and reconstructed data, and can be reduced by increasing marginal likelihood according to Equation (9). In the context of IRT, AVB thus stands out by capturing a more accurate respondents’ latent traits. Research in the computer science field also highlights that the combination of GAN and VAE can extract more discriminative information and provide more precise and clear model in the latent variable space (e.g., Makhzani et al.,, 2015; Hou et al.,, 2019).
@@ 3.1 Neural Network Design
Parameter initialization is a critical step in training neural networks, as it sets the initial values of weights and biases before optimization begins. Proper initialization prevents issues like vanishing or exploding gradients, which can hinder learning and slow convergence. For encoder network and discriminator network, Kaiming initialization (He et al.,, 2015) is chosen for the weight and bias since it is tailored for asymmetric activation function like GELU. As GELU almost drops most negative activations and reduces variance of parameters in subsequent layers, Kaiming initialization preserve variance across layers. Considering $P_{l}\times P_{l-1}$ matrix $\bm{W}_{0}^{(l)}=[w_{p_{l},p_{l-1},0}]$ and $P_{l}\times 1$ vector $\bm{b}_{0}^{(l)}=(b_{p_{l},0})$, Kaiming initialization is defined as follows:

### v2

@@ Abstract
Advances in deep learning and representation learning have transformed item factor analysis (IFA) in the item response theory (IRT) literature by enabling more efficient and accurate parameter estimation. Variational Autoencoders (VAEs) have been one of the most impactful techniques in modeling high-dimensional latent variables in this context. However, the limited expressiveness of the inference model based on traditional VAEs can still hinder the estimation performance. We introduce Adversarial Variational Bayes (AVB) algorithms as an improvement to VAEs for IFA with improved flexibility and accuracy. By bridging the strengths of VAEs and Generative Adversarial Networks (GANs), AVB incorporates an auxiliary discriminator network to reframe the estimation process as a two-player adversarial game and removes the restrictive assumption of standard normal distributions in the inference model. Theoretically, AVB can achieve similar or higher likelihood compared to VAEs. A further enhanced algorithm, Importance-weighted Adversarial Variational Bayes (IWAVB) is proposed and compared with Importance-weighted Autoencoders (IWAE). In an exploratory analysis of empirical data, IWAVB demonstrated superior expressiveness by achieving a higher likelihood compared to IWAE. In confirmatory analysis with simulated data, IWAVB achieved similar mean-square error results to IWAE while consistently achieving higher likelihoods. When latent variables followed a multimodal distribution, IWAVB outperformed IWAE. With its innovative use of GANs, IWAVB is shown to have the potential to extend IFA to handle large-scale data, facilitating the potential integration of psychometrics and multimodal data analysis.
@@ 1 Introduction
Psychological and educational research now frequently involves diverse data types, including large-scale assessment, high-volume data, and process data involving unstructured and textual responses (e.g., Weston and Laursen,, 2015; Zhang et al.,, 2023; Ma et al.,, 2024). Traditional estimation algorithms often struggle to efficiently process and analyze these data, leading to unstable model estimates, long computation times, and excessive memory usage. Deep Learning (DL) literature has much to offer to handle the above challenges. If we view IFA from the lens of generalized latent variable models (Muthén,, 2002; Rabe-Hesketh and Skrondal,, 2008), representation learning (Bengio et al.,, 2013) can help model and capture intricate latent patterns from observed data. Unlike traditional IFA estimation methods, which often rely on strict parametric and distributional assumptions, representation learning leverages powerful learning algorithms and expressive models such as neural networks to generate richer and highly flexible representations when modeling data. This improves the modeling and detection of subtle patterns for high-dimensional latent variables that classical methods might underfit. Therefore, by integrating representation learning with IFA, researchers can achieve a more comprehensive understanding of latent traits.
@@
An alternative representation paradigm, Generative Adversarial Networks (GANs, Goodfellow et al.,, 2014), a celebrated framework for estimating generative models in deep learning and computer vision, offer greater flexibility in modeling arbitrary probability densities and can contribute to the expressiveness. Multiple variants of GANs have achieved remarkable success in complex tasks such as image generation (e.g., Radford,, 2015; Arjovsky et al.,, 2017; Zhu et al., 2017a, ). Among these models, the combination of GANs and VAEs can be traced back to Adversarial Autoencoders (AAEs; Makhzani et al.,, 2015) and the VAE-GAN model (Larsen et al.,, 2016). Figure 1 provides an intuitive visual example demonstrating how integrating GANs and VAEs can lead to better representation learning. The figure compares ten handwritten images sampled from four models: (a) GAN, (b) WGAN (a GAN variant, Arjovsky et al.,, 2017), (c) VAE, and (d) VAE-GAN. While GAN generated only the digit 1, WGAN exhibited unstable performance, and the images produced by VAE appeared blurry. In contrast, VAE-GAN generated relatively clear results. However, vanilla GANs, VAE-GANs, and AAEs rely solely on adversarial loss, formulated as a minimax game between networks to deceive a discriminator. As a result, they are not motivated from a maximum-likelihood perspective and are therefore not directly applicable for improving Marginal Maximum Likelihood (MML) estimation in IFA.
@@ 2 Adversarial Variational Bayes for Item Factor Analysis
In the context of IFA, given the observed responses $\bm{x}$, the Variational Inference (VI) method uses an inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ to approximate the posterior distribution of the latent variable $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$. The $\bm{\phi}$ represents the parameters for the inference model, and ${\bm{\theta}}$ includes the parameters for the posterior distributions (e.g., items’ parameters). As a specific approach to perform VI, Adversarial Variational Bayes can be regraded as an improved version of Variational Autoencoder which constructs more expressive inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$, and is shown to improve the likelihood estimation of parameters for the Graded Response Model in this section.
@@ 2.1 The Graded Response Model
The Graded Response Model (GRM; Samejima,, 1969) is a specific model for IFA used for analyzing ordinal response data, particularly in psychological and educational measurement. The GRM assumes responses to an item are ordered categories (e.g., Likert scale). We follow the notation as in Cai, (2010).
Given $i=1,\dots,N$ distinct respondents and $j=1,\dots,M$ items, the response for respondent $i$ to item $j$ in $C_{j}$ graded categories is $x_{i,j}\in\left\{0,1,\dots,C_{j}-1\right\}$. Assuming the number of latent variables is $P$, the $P\times 1$ vector of latent variables for respondent $i$ is denoted as $\bm{z}_{i}$. For item $j$, the $P\times 1$ vector of loadings is $\bm{\beta}_{j}$, and the $(C_{j}-1)\times 1$ vector of strictly ordered category intercepts is $\bm{\alpha}_{j}=(\alpha_{j,1},\dots,\alpha_{j,C_{j}-1})^{T}$. Also, the parameters of item $j$ are denoted as $\bm{\theta}_{j}=(\bm{\alpha}_{j},\bm{\beta}_{j})$. The Graded Response Model is defined as the boundary probability conditional on the item parameters $\bm{\theta}_{j}$ and latent variables $\bm{z}_{i}$: for $k\in\left\{1,\dots,C_{j}-1\right\}$,
@@
The conditional probability for a specific response $x_{i,j}=k\in\left\{0,\dots,C_{j}-1\right\}$ is
@@ 2.2 Marginal Maximum Likelihood
For the respondent $i$, and the item $j$, with the Equation (2), the response $x_{i,j}$ follows the multinomial distribution with 1 trail and $C_{j}$ mutually exclusive outcomes. The probability mass function is: $p_{\bm{\theta_{j}}}(x_{i,j}\mid\bm{z}_{i})=\prod_{k=0}^{C_{j}-1}p_{i,j,k}^{\mathbbm{1}_{k}(x_{i,j})}$, where $\mathbbm{1}_{k}(x_{i,j})=1$ if $x_{i,j}=k$ and otherwise it is 0.
Considering all the responses $\bm{x}_{i}=(x_{i,j})_{j=1}^{M}$ of the respondent $i$, since each response is conditionally independent given the latent variables, the conditional probability is: for $\bm{\theta}=(\bm{\theta}_{j})_{j=1}^{M}$,
$p_{\bm{\theta}}(\bm{x}_{i}\mid\bm{z}_{i})=\prod\limits_{j=1}^{M}p_{\bm{\theta_{j}}}(x_{i,j}\mid\bm{z}_{i}).$ (3)
Given the prior distribution of latent variables is $p(\bm{z}_{i})$ and $p_{\bm{\theta}}(\bm{x}_{i}\mid\bm{z}_{i})$, the marginal likelihood of $\bm{x}_{i}$ is:
@@
Since the above equation contains $N$ integrals in high-dimensional latent space $\mathbb{R}^{P}$, directly maximizing the likelihood is computationally intensive. Instead, variational methods that the current study focuses on improving, approximate $\log\mathcal{L}(\bm{\theta}\mid\bm{X})$ by one lower bound and achieve efficient computation.
@@ 2.3 Variational Inference and Variational Autoencoder
For each set of item parameters $\bm{\theta}$, if there exists $\bm{\phi}$ such that $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ exactly match $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$ the KL divergence term will be zero, so the MML estimator $\bm{\theta}^{*}$ will also be the maximizer of ELBO. However, this will be a very ideal case. Generally speaking, for fixed $\bm{\theta}$, if ELBO is increased by taking gradient steps in $\bm{\phi}$ space, the KL divergence will more approach to its minimum. On the other hand, if $\bm{\phi}$ is fixed, increasing ELBO with respect to $\bm{\theta}$ will push the marginal likelihood $p_{\bm{\theta}}(\bm{x})$ higher. Therefore, our objective is
@@ 2.4 Adversarial Variational Bayes
Therefore, we need to do two optimizations for Objectives (11) and (13), summarized in Algorithm 1. Note that $T(\bm{x},\bm{z})$ and $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ can be implemented as neural networks, but in the context of the GRM, the decoder $p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)$ corresponds directly to the likelihood definition for the GRM. As such, it is modeled by a single-layer neural network consisting of a linear function and a sigmoid activation.
@@ 2.5 Our Contributions to the Literature
Given that VAEs have been a powerful approach to perform exploratory IFA (Urban and Bauer,, 2021), the innovative contribution of the AVB method is to bring richer inference model choices for the implementation by deep learning. Over the past decade, deep learning techniques have thrived (LeCun et al.,, 2015), and IFA has derived considerable benefits from these advances. When applying neural networks to estimate parameters in IFA, key distinctions arise in the design of the network architecture and the choice of loss function. A straightforward strategy is to feed one-hot encodings [^1] In one-hot encoding, each respondent or item is represented by a binary vector whose length matches the total number of respondents or items. Exactly one element of this vector is set to 1 (indicating the specific respondent or item), and all other elements are set to 0. for both respondents and items into two separate feedforward neural networks: one infers the latent variables of the respondents, while the other estimates the item parameters. This approach typically employs a variation of the Joint Maximum Likelihood (JML) for its loss function, relying on the product of conditional probabilities in Equation (3) across all respondents (Tsutsumi et al.,, 2021; Zhang and Chen,, 2024). However, as demonstrated by Ghosh, (1995), the JML solution leads to inconsistent estimates for item parameters.
An alternative that proves more statistically robust relies on the marginal maximum likelihood procedure discussed earlier (Bock and Aitkin,, 1981; Wirth and Edwards,, 2007). Drawing on these perspectives, researchers merge representation learning for latent-variable models with neural architectures by adopting autoencoder designs. Through variational methods, they approximate maximum likelihood estimation for latent variable models. A notable example in the literature is the Importance-Weighted Autoencoder algorithm (Urban and Bauer,, 2021). In our study, we propose an importance-weighted adversarial variational Bayes, refines this approach.
@@
To address this limitation, the AVB framework modifies $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ from Equation (10) into a fully black-box inference mechanism that cooperates with adversarial objectives. Through this process, the algorithm derives item parameter estimates $\bm{\theta}^{*}$ that optimizes the likelihood of the observed data, while also constructing a distribution $q_{\bm{\phi}^{*}}(\bm{z}\mid\bm{x})$ that closely approximates the true posterior $p_{\bm{\theta}^{*}}(\bm{z}\mid\bm{x})$.
@@
Under misspecification such as inappropriate assumptions of latent variable distributions, AVB’s architecture can deliver a higher marginal likelihood theoretically. Meanwhile, improving the marginal log-likelihood can reduce the reconstruction error $-\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]$, which measures the disparity between real and reconstructed data, and can be reduced by increasing marginal likelihood according to Equation (8). In the context of IRT, AVB thus stands out by capturing a more accurate respondents’ latent traits. Research in the computer science field also highlights that the combination of GAN and VAE can extract more discriminative information and provide more precise and clear model in the latent variable space (e.g., Makhzani et al.,, 2015; Hou et al.,, 2019).
@@ 3.1 Neural Network Design
Parameter initialization is a critical step in training neural networks, as it sets the initial values of weights and biases before optimization begins. Proper initialization prevents issues like vanishing or exploding gradients, which can hinder learning and slow convergence. For encoder network and discriminator network, Kaiming initialization (He et al.,, 2015) is chosen for the weight and bias since it is tailored for asymmetric activation function like GELU. As GELU almost drops most negative activations and reduces variance of parameters in subsequent layers, Kaiming initialization preserves variance across layers. Considering $P_{l}\times P_{l-1}$ matrix $\bm{W}_{0}^{(l)}=[w_{p_{l},p_{l-1},0}]$ and $P_{l}\times 1$ vector $\bm{b}_{0}^{(l)}=(b_{p_{l},0})$, Kaiming initialization is defined as follows: